OpenStudy (anonymous):

solve for x x^2+15x+36=0

OpenStudy (anonymous):

(x+3)(x+12)=0

OpenStudy (anonymous):

x=-3 x=-12

OpenStudy (anonymous):

@elisasaab I plug them back in and they don't equal to zero

OpenStudy (anonymous):

@bjschultze87 read on for x=-3 and x =-12 \[ x^2+15x+36 (-3)^2 + 15(-3) +36= 9-45 +36=0\\ (-12)^2 + 15(-12) +36=144-180 +36=0 \]